# GenMix: Effective Data Augmentation with Generative Diffusion Model Image Editing

### v3

@@ 4 Experimental Setup
Datasets. To enable comparisons with previous research on image augmentation Yun et al. (2019b); Zhang et al. (2018); Kang and Kim (2023); Uddin et al. (2020); Hendrycks et al. (2019); Cheung and Yeung (2021); Cubuk et al. (2020, 2019); Kim et al. (2020a, b); Huang et al. (2021); Lim et al. (2019); Verma et al. (2019), we assess our approach using both general image classification and fine-grained image classification datasets. For general image classification, we use three datasets: ImageNet-1K Deng et al. (2009), CIFAR100 Krizhevsky et al. (2009), and Tiny-ImageNet-200 Le and Yang (2015). For fine-grained image classification, we employ three datasets: CUB Nilsback and Zisserman (2008), Stanford Cars Krause et al. (2013), and Aircraft Maji et al. (2013). Moreover, we use Flower102 Nilsback and Zisserman (2008), Aircraft Maji et al. (2013), and Stanford Cars Krause et al. (2013) for evaluation related to Transfer Learning. The data scarcity related experiments are conducted with Flower102 Nilsback and Zisserman (2008) dataset. Our experiments for domain adaption additionally use OfficeHome Venkateswara et al. (2017) dataset. These datasets encompass a variety of scenarios, featuring images with a broad spectrum of objects like plants, animals, vehicles, human actions, and general objects across diverse scenes and textures.
@@
Implementation Details. For image editing, we employed InstructPix2Pix pre-trained image editing diffusion model Brooks et al. (2023) utilizing our custom-built textual prompts. To create the Mask $M_{u}$ we used a fixed blending width of 20 pixels. We use $\lambda=0.20$ for blending the fractal image in Eq. (4).
@@
Textual Prompt Selection. To ensure that only suitable prompts are used for classification, a custom textual library of filter-like global visual effects is predefined, comprising ‘autumn’, ‘snowy’, ‘sunset’, ‘watercolor art’, ‘rainbow’, ‘aurora’, ‘mosaic’, ‘ukiyo-e’ and ‘a sketch with crayon’. For domain adaptation, we introduce an additional set of bespoke conditional prompts, such as ‘graffiti’, ‘retro comic’, ‘chalk drawing’, ‘watercolor painting’, ‘digital art’ and ‘cartoon style’. These prompts are designed to introduce a broader range of visual appearances, enhancing the model’s adaptability to different domains ensuring their generic nature and wide applicability. These prompts do not significantly edit the image structure while introducing global visual effects. Each prompt in the textual library is appended with the template ‘A transformed version of image into $prompt$’ to form a specific input for the image editing diffusion model. Examples of images generated using these prompts are shown in Figure 3.
TABLE II: Top-1 and Top-5 accuracies comparison on ImageNet-1K using ResNet-50, on Tiny ImageNet-200 and CIFAR-100 using PreActResNet-18. FGSM error rates on CIFAR-100 and Tiny-ImageNet-200 datasets are also computed for PreActResNet-18. Compared numbers are taken either from the original papers or from Kang and Kim (2023), following the exact protocols.
Method ImageNet-1K Tiny ImageNet-200 CIFAR-100
Top-1 Acc (%) Top-5 Acc (%) Top-1 Acc (%) Top-5 Acc (%) FGSM Error (%) Top-1 Acc (%) Top-5 Acc (%) FGSM Error (%)
Vanilla-He et al. (2016b) 75.97 92.66 57.23 73.65 42.77 76.33 91.02 23.67
AugMix Hendrycks et al. (2019) 76.75 93.30 55.97 74.68 - 75.31 91.62 43.33
ManifoldMix Verma et al. (2019) 76.85 93.50 58.01 74.12 41.99 79.02 93.37 20.98
Mixup Zhang et al. (2018) 77.03 93.52 56.59 73.02 43.41 76.84 92.42 23.16
CutMix Yun et al. (2019b) 77.08 93.45 56.67 75.52 43.33 76.80 91.91 23.20
Guided-SR Kim et al. (2020a) 77.20 93.66 55.97 74.68 - 80.60 94.00 -
PixMix Hendrycks et al. (2022) 77.40 - - - - 79.70 - -
PuzzleMix Kim et al. (2020a) 77.51 93.76 63.48 75.52 36.52 80.38 94.15 19.62
GuidedMix Kang and Kim (2023) 77.53 93.86 64.63 82.49 - 81.20 94.88 -
Co-Mixup Kim et al. (2020b) 77.63 93.84 64.15 - - 80.15 - -
YOCO Han et al. (2022) 77.88 - - - - - - -
Azizi et al. Azizi et al. (2023) 78.17 - - - - - - -
GMix 78.73 95.47 65.80 83.70 34.47 82.58 95.51 16.83
5 Performance Evaluation
5.1 General Classification
@@
General Classification (GC) is an essential task used to assess the effectiveness of data augmentation approaches. An increase in GC accuracy indicates that the augmentation method fairly introduces meaningful variations in input data to enhance the learning process. In our evaluation, we applied our method to three challenging datasets: CIFAR-100, ImageNet-1K, and Tiny-ImageNet-200. Aligning with the practices of current state-of-the-art (SOTA) methods, we used ResNet50 for the ImageNet-1K, and utilized the PreActResNet18 for the Tiny-ImageNet-200 and CIFAR-100 datasets Huang et al. (2021); Kang and Kim (2023); Kim et al. (2020b); Uddin et al. (2020), and employed ResNet50 for the ImageNet dataset Kim et al. (2020a); Hendrycks et al. (2019); Yun et al. (2019b).
@@
Table II showcases a comparison of Top-1 and Top-5 performances on ImageNet-1K dataset, and compares with the existing SOTA methods. GMix exhibited strong performance, achieving a Top-1 accuracy gain of $2.76\%$ over the Vanilla He et al. (2016b) performer. Compared to GuidedMixup Kang and Kim (2023), which shows the second best results for Top-5 accuracy, our approach attains an absolute performance improvement of $1.6\%$ for Top-5 and $1.2\%$ for Top-1 accuracy. Table II also showcases a comparison of performances on Tiny-ImageNet-200 with existing SOTA methods. Our proposed GMix demonstrates superior performance over existing data augmentation approaches. Over the Vanilla He et al. (2016b), we obtained absolute Top-1 and Top-5 performance gains of $8.57\%$ and $10.05\%$, respectively. We outperform the strong baseline of GuidedMixup Kang and Kim (2023) for Top-5 accuracy by an absolute gain of $1.2\%$.
@@
We observe similar trends on the CIFAR-100 dataset. GMix outperforms the Vanilla He et al. (2016b) with Top-1 and Top-5 absolute accuracy gains of $6.25\%$ and $4.49\%$, respectively. These results across diverse and challenging benchmark datasets underscore the effectiveness of GMix in enhancing learning outcomes. They also indicate its capability to mitigate overfitting and promote better generalization in neural network training.
5.2 Adversarial Robustness
@@
In line with current state-of-the-art methods Kim et al. (2020a); Yun et al. (2019b); Hendrycks et al. (2019); Verma et al. (2019); Uddin et al. (2020); Hong et al. (2021); Kim et al. (2020b); Kang and Kim (2023), we assess the robustness of GMix method against adversarial attacks and input perturbations for general classification. For these evaluations, we employ fast adversarial training Wong et al. (2020) to generate adversarially perturbed input images. The primary objective of these tests is to determine if our augmentation technique offers enhanced resistance to adversarial attacks. We measure performance using the FGSM Wong et al. (2020) error rates as an indicator of robustness against such attacks.
@@
As presented in Table II, our approach achieves an error rate of $16.83\%$ on CIFAR-100, outperforming all other methods in comparison. Similarly, on Tiny ImageNet-200, our method outperforms existing state-of-the-art techniques with an error rate of $34.47\%$. These results highlight that our method maintains strong resilience to adversarial perturbations, surpassing the performance of existing top-performing approaches.
5.3 Fine-Grained Visual Classification
@@
Fine-Grained Visual Classification (FGVC) poses arduous challenge compared to general classification because it requires distinguishing subtle variations between objects within the similar category with broader range. To effectively train a model on such tasks, a generative augmentation approach must carefully preserve these fine-grained details. We assess the performance of GMix on FGVC task using three well-known datasets: CUB Nilsback and Zisserman (2008), Stanford Cars Krause et al. (2013), and Aircraft Maji et al. (2013), employing the ResNet-50 He et al. (2016b) architecture. As demonstrated in Table III, GMix outperforms other competitive methods, showing its strength in enhancing the model’s ability to generalize.
@@
It is notable that GMix achieves the highest accuracy across all the datasets. On the CUB dataset, it achieves an absolute gain of top $2.3\%$ over the strong baseline of Guided-AP Kang and Kim (2023). For the Aircraft dataset, GMix achieves an absolute gain of $1.5\%$. Lastly, on the Stanford Cars dataset, the absolute gain is again more than 1%. This consistent performance gain of GMixcan be attributed to the preservation of subtle but important visual information in the hybrid image constructed in our approach.
TABLE III: Top-1 (%) performance comparison on fine-grained task of ResNet-50.
Method Birds Aircraft Cars
Vanilla  He et al. (2016b) 65.50 80.29 85.52
Simple He et al. (2016b) - 80.13 86.36
Auto Augment Cubuk et al. (2019) - 82.28 88.04
Fast AutoAugment Lim et al. (2019) - 82.20 87.19
DADA Li et al. (2020) - 81.16 87.14
RandAugment Cubuk et al. (2020) - 82.30 87.79
AdaAug Cheung and Yeung (2021) - 82.50 88.49
Mixup Zhang et al. (2018) 71.33 82.38 88.14
CutMix Yun et al. (2019b) 72.58 82.45 89.22
SaliencyMix Uddin et al. (2020) 66.66 83.14 89.04
Guided-SR Kang and Kim (2023) 74.08 83.51 89.23
SnapMix  Huang et al. (2021) 75.53 82.96 90.10
PuzzleMix  Kim et al. (2020a) 74.85 82.66 89.68
Co-Mixup  Kim et al. (2020b) 72.83 83.57 89.53
Guided-AP  Kang and Kim (2023) 77.08 84.32 90.27
GMix 79.42 85.84 91.30
TABLE IV: Top-1 (%) accuracy of GMix on fine-tuning experiments using ImageNet pretrained ResNet-50.
Method Flower102 Aircraft Cars
Vanilla He et al. (2016b) 94.98 81.60 88.08
AA Cubuk et al. (2019) 93.88 83.39 90.82
RA Cubuk et al. (2020) 95.23 82.98 89.28
Fast AA Lim et al. (2019) 96.08 82.56 89.71
AdaAug Cheung and Yeung (2021) 97.19 83.97 91.18
GMix 98.27 85.83 93.47
5.4 Transfer Learning
@@
Transfer learning or fine-tuning, is a prevalent technique for adapting large neural network architectures pre-trained on large-scale datasets to smaller datasets. It requires limited computational resources and helps conducting experiments rapidly. Notably, many augmentation methods Yun et al. (2019b); Zhang et al. (2018); Kang and Kim (2023); Uddin et al. (2020); Hendrycks et al. (2019); Kim et al. (2020a, b); Huang et al. (2021); Verma et al. (2019) have not reported their performance for this popular application. Nevertheless, we evaluate our approach by fine-tuning the baseline model on three datasets Flower102, Aircraft, and Stanford Cars, utilizing ResNet-50 model pre-trained on ImageNet. The results are summarized in Table IV.
@@
On the Flower102 dataset, GMix achieved the highest accuracy of $98.27\%$. A similar trend is evident in the Aircraft dataset, where GMix reached $85.83\%$. In the Stanford Cars dataset, GMix again led with $93.47\%$. GMix significantly surpassed other augmentation techniques, demonstrating its effectiveness in transfer learning scenarios. As compared to the vanilla baselines, our method achieves absolute performance gains of $3.3\%$, $4.2\%$ and $5.4\%$ on the used datasets. It is noteworthy that the accuracy obtained by GMix through fine-tuning (Table IV) is comparable to its performance when trained from scratch on the Aircraft and Cars dataset (Table III). Since fine-tuning requires substantially fewer computational resources than training from scratch, this experiment highlights the practical importance of GMix in deep learning applications.
5.5 Data Scarcity
@@
Data scarcity presents a significant challenge when training deep neural networks. With only a limited number of examples per class, deep networks may struggle to learn meaningful patterns, which can lead to overfitting and a reduced ability to generalize. To mitigate this, augmentation techniques are commonly applied to increase the amount of training data. In this context, we evaluate the performance of ResNet-18 He et al. (2016b) trained with just $10$ images per class from the original Flower102 dataset. As shown in Table V, GMix consistently outperforms other mixup approaches, achieving Top-1 accuracy of $77.23\%$ and Top-5 accuracy of $74.35\%$. Our method is specifically designed to diversify the training dataset. By using custom conditional prompts, GMix artificially expands and enriches the dataset, leading to stronger neural network learning.
TABLE V: Top-1 (%) accuracy of ResNet-18 on the data scarcity task using the Flower102 dataset, where only 10 images per class are present in the dataset.
Method Valid Test
Vanilla He et al. (2016b) 64.48 59.14
Mixup Zhang et al. (2018) 70.55 66.81
CutMix Yun et al. (2019b) 62.68 58.51
SaliencyMix Uddin et al. (2020) 63.23 57.45
Guided-SR Kang and Kim (2023) 72.84 69.31
SnapMix Huang et al. (2021) 65.71 59.79
PuzzleMix Kim et al. (2020a) 71.56 66.71
Co-Mixup Kim et al. (2020b) 68.17 63.20
GuidedMixup Kang and Kim (2023) 74.74 70.44
GMix 77.23 74.35
@@
TABLE VI: Comparison of H-score (%) in the Open Partial Domain Adaptation (OPDA) scenario on the Office-Home dataset, where Clipart (Cl), Art (Ar), Real (Re), and Product (Pr) are different domains. Unified (U) methods are applicable for all scenarios including OPDA, OSDA, and PDA. The SF denotes source data-free, and the CF indicates K-means clustering free model adaption. Our method is shown as +GMix to imply LEAD+GMix. Boldface represents performance gains of GMixover LEAD.
Methods U SF CF Ar2Cl Ar2Pr Ar2Re Cl2Ar Cl2Pr Cl2Re Pr2Ar Pr2Cl Pr2Re Re2Ar Re2Cl Re2Pr Avg
CMU Fu et al. (2020) 56.0 56.9 59.2 67.0 64.3 67.8 54.7 51.1 66.4 68.2 57.9 69.7 61.6
DANCE Saito et al. (2020a) 61.0 60.4 64.9 65.7 58.8 61.8 73.1 61.2 66.6 67.7 62.4 63.7 63.9
DCC Li et al. (2021) 58.0 54.1 58.0 74.6 70.6 77.5 64.3 73.6 74.9 81.0 75.1 80.4 70.2
OVANet Saito and Saenko (2021a) 62.8 75.6 78.6 70.7 68.8 75.0 71.3 58.6 80.5 76.1 64.1 78.9 71.8
GATE Chen et al. (2022b) 63.8 75.9 81.4 74.0 72.1 79.8 74.7 70.3 82.7 79.1 71.5 81.7 75.6
UniOT Chang et al. (2022) 67.3 80.5 86.0 73.5 77.3 84.3 75.5 63.3 86.0 77.8 65.4 81.9 76.6
@@
SHOT-O Liang et al. (2020) 32.9 29.5 39.6 56.8 30.1 41.1 54.9 35.4 42.3 58.5 33.5 33.3 40.7
LEAD Qu et al. (2024) 62.7 78.1 86.4 70.6 76.3 83.4 75.3 60.6 86.2 75.4 60.7 83.7 75.0
+ GMix 63.9 83.0 85.9 76.8 78.9 85.3 78.1 62.2 86.9 79.8 62.9 83.0 77.2
TABLE VII: H-score (%) comparison in the Open-set Domain Adaptation (OSDA) scenario on the Office-Home dataset.
Methods U SF CF Ar2Cl Ar2Pr Ar2Re Cl2Ar Cl2Pr Cl2Re Pr2Ar Pr2Cl Pr2Re Re2Ar Re2Cl Re2Pr Avg
CMU Fu et al. (2020) 55.0 57.0 59.0 59.3 58.2 60.6 59.2 51.3 61.2 61.9 53.5 55.3 57.6
DANCE Saito et al. (2020a) 6.5 9.0 9.9 20.4 10.4 9.2 28.4 12.8 12.6 14.2 7.9 13.2 12.9
DCC Li et al. (2021) 56.1 67.5 66.7 49.6 66.5 64.0 55.8 53.0 70.5 61.6 57.2 71.9 61.7
OVANet Saito and Saenko (2021a) 58.6 66.3 69.9 62.0 65.2 68.6 59.8 53.4 69.3 68.7 59.6 66.7 64.0
@@
SHOT-O Liang et al. (2020) 37.7 41.8 48.4 56.4 39.8 40.9 60.0 41.5 49.7 61.8 41.4 43.6 46.9
LEAD Qu et al. (2024) 60.7 70.8 76.5 61.0 68.6 70.8 65.5 59.8 74.2 64.8 57.7 75.8 67.2
+ GMix 62.8 72.4 76.8 65.1 66.8 71.5 67.3 58.5 74.9 65.7 63.1 79.2 68.7
5.6 Domain Adaptation

### v4

@@ IV Experimental Setup
Datasets. To enable comparisons with previous research on image augmentation [27, 28, 24, 22, 37, 64, 63, 68, 19, 21, 20, 69, 60], we assess our approach using both general image classification and fine-grained image classification datasets. For general image classification, we use three datasets: ImageNet-1K [70], CIFAR100 [71], and Tiny-ImageNet-200 [72]. For fine-grained image classification, we employ three datasets: CUB [73], Stanford Cars [74], and Aircraft [75]. Moreover, we use Flower102 [73], Aircraft [75], and Stanford Cars [74] for evaluation related to Transfer Learning. The data scarcity related experiments are conducted with Flower102 [73] dataset. Our experiments for domain adaption additionally use OfficeHome [76] dataset. These datasets encompass a variety of scenarios, featuring images with a broad spectrum of objects like plants, animals, vehicles, human actions, and general objects across diverse scenes and textures.
@@
Implementation Details. For image editing, we employed InstructPix2Pix pre-trained image editing diffusion model [77] utilizing our custom-built textual prompts. To create the Mask $M_{u}$ we used a fixed blending width of 20 pixels. We use $\lambda=0.20$ for blending the fractal image in Eq. (5).
@@
Textual Prompt Selection. To ensure that only suitable prompts are used for classification, a custom textual library of filter-like global visual effects is predefined, comprising ‘autumn’, ‘snowy’, ‘sunset’, ‘watercolor art’, ‘rainbow’, ‘aurora’, ‘mosaic’, ‘ukiyo-e’ and ‘a sketch with crayon’. For domain adaptation, we introduce an additional set of bespoke conditional prompts, such as ‘graffiti’, ‘retro comic’, ‘chalk drawing’, ‘watercolor painting’, ‘digital art’ and ‘cartoon style’. These prompts are designed to introduce a broader range of visual appearances, enhancing the model’s adaptability to different domains ensuring their generic nature and wide applicability. These prompts do not significantly edit the image structure while introducing global visual effects. Each prompt in the textual library is appended with the template ‘A transformed version of image into promptprompt’ to form a specific input for the image editing diffusion model. Examples of images generated using these prompts are shown in Figure 4.
TABLE III: FGSM error rates on CIFAR-100 and Tiny-ImageNet-200 datasets for PreActResNet-18, following the protocols previous SOTA mixup methods [53, 24] and we reproduce results for AdAutoMix [30].
Dataset Vanilla Mixup Manifold CutMix AugMix PuzzleMix AdAutoMix DiffuseMix GenMix
CIFAR-100 23.67 23.16 20.98 23.20 43.33 19.62 24.34 17.38 16.83
Tiny-ImageNet-200 42.77 43.41 41.99 43.33 36.52 34.53 33.83
V Performance Evaluation
V-A General Classification
@@
General Classification (GC) is an essential task used to assess the effectiveness of data augmentation approaches. An increase in GC accuracy indicates that the augmentation method fairly introduces meaningful variations in input data to enhance the learning process. In our evaluation, we applied our method to three challenging datasets: CIFAR-100 [71], ImageNet-1K [70], and Tiny-ImageNet-200 [72]. Aligning with the practices of current state-of-the-art (SOTA) methods, we used ResNet-50 for the ImageNet-1K, and utilized ResNet-18 [2] and ResNeXt-50 [78] for the Tiny-ImageNet-200 and CIFAR-100 datasets [20, 24, 21, 22], and employed ResNet-50 for the ImageNet dataset [19, 37, 27].
@@
Table II showcases a comparison of Top-1 performances on the ImageNet-1K dataset and compares with the existing SOTA methods. GenMix exhibited strong performance, achieving a Top-1 accuracy gain of $2.49\%$ over the Vanilla [66] baseline. Compared to DiffuseMix [53], which shows the second-best Top-1 result of $78.64\%$, our approach attains an absolute performance improvement of $0.68\%$. Table II also showcases a comparison of performances on Tiny-ImageNet-200 with existing SOTA methods. Our proposed GenMix demonstrates superior performance over existing data augmentation approaches. Over the Vanilla [66], we obtained an absolute Top-1 performance gain of $9.77\%$, and an improvement of $1.19\%$ over DiffuseMix [53].
@@
We observe similar trends on the CIFAR-100 dataset. GenMix outperforms the Vanilla [66] with a Top-1 absolute accuracy gain of $6.21\%$ using ResNet-18. Compared to DiffuseMix [53], our method achieves a gain of $1.47\%$ on the same backbone. These results across diverse and challenging benchmark datasets underscore the effectiveness of GenMix in enhancing learning outcomes. They also indicate its capability to mitigate overfitting and promote better generalization in neural network training.
V-B Adversarial Robustness
@@
In line with current state-of-the-art methods [19, 27, 37, 60, 22, 79, 21, 24, 80], we assess the robustness of GenMix method against adversarial attacks and input perturbations for general classification. For these evaluations, we employ fast adversarial training [81] to generate adversarially perturbed input images. The primary objective of these tests is to determine if our augmentation technique offers enhanced resistance to adversarial attacks. We measure performance using the FGSM [81] error rates as an indicator of robustness against such attacks.
@@
As presented in Table III, our approach achieves an FGSM error rate of $16.83\%$ on CIFAR-100, outperforming all other methods in comparison. Similarly, on Tiny ImageNet-200, our method outperforms existing state-of-the-art techniques with an error rate of $33.83\%$. These results highlight that our method maintains strong resilience to adversarial perturbations, surpassing the performance of existing top-performing approaches.
TABLE IV: Top-1 (%) performance comparison for fined-grained tasks. Previous results are taken from AdAutoMix [30] and reproduce DiffuseMix [53] results on new backbones.
Method FGVC-Aircrafts Stanford-Cars
ResNet18 ResNeXt50 ResNet18 ResNeXt50
Vanilla [66] 80.23 85.10 86.32 90.15
Mixup [28] 79.52 85.18 86.27 90.81
CutMix [27] 78.84 84.55 87.48 91.22
ManiMix [60] 80.68 86.60 85.88 90.20
SalieMix [22] 80.02 84.31 85.88 90.20
FMix [36] 79.36 86.23 87.55 90.90
PuzzleMix [20] 80.76 86.23 87.78 91.29
ResizeMix [19] 78.10 84.08 88.17 91.36
AutoMix [21] 81.37 86.72 88.89 91.38
AdAuto[24] 81.73 87.16 89.19 91.59
DiffuseMix[53] 82.69 88.13 90.25 92.36
GenMix 83.87 89.27 91.54 93.21
TABLE V: Top-1 (%) accuracy on the data scarcity experiment under the setting ($10$ images per class).
Method Vanilla Mixup CutMix SaliencyMix SnapMix PuzzleMix Co-Mixup GuidedMixup DiffuseMix GenMix
Validation 64.48 70.55 62.68 63.23 65.71 71.56 68.17 74.74 77.14 79.34
Test 59.14 66.81 58.51 57.45 59.79 66.71 63.20 70.44 74.12 76.38
TABLE VI: Top-1 (%) performance on ResNet50 with CUB200 and Standford-Cars datasets.
Method Vanilla Mixup CutMix PuzzleMix AutoMix AdAutoMix DiffuseMix GenMix
CUB-Birds 200 81.76 82.79 81.67 82.59 82.93 83.36 83.83 84.45
Stanford-Cars 88.88 89.45 88.99 89.37 88.71 89.65 90.26 90.84
V-C Fine-Grained Visual Classification
@@
Fine-Grained Visual Classification (FGVC) poses arduous challenge compared to general classification because it requires distinguishing subtle variations between objects within the similar category with broader range [82]. To effectively train a model on such tasks, a generative augmentation approach must carefully preserve these fine-grained details. We assess the performance of GenMix on FGVC task using two well-known datasets: Stanford Cars [74], and FGVC-Aircraft [75], employing the ResNet-50 [66] and ResNeXt [83] architectures.
@@
As demonstrated in Table IV, GenMix outperforms other competitive methods, showing its strength in enhancing the model’s ability to generalize. It is notable that GenMix achieves the highest accuracy across all the datasets. On the FGVC-Aircraft dataset, it achieves an absolute gain of top $1.18\%$ on ResNet18 and $1.14\%$ on ResNext50 over the strong baseline of DiffuseMix  [53]. For the Stanford-Cars dataset, GenMix achieves an absolute gain of $1.29\%$ on ResNet18 and $0.85\%$ on ResNext50. Lastly, on the Stanford Cars dataset, the absolute gain is again more than  1%. This consistent performance gain of GenMix can be attributed to the preservation of subtle visual information in the hybrid image constructed in our approach.
V-D Transfer Learning
@@
Transfer learning is a prevalent technique for adapting large neural network architectures pre-trained on large-scale datasets to smaller datasets. It requires limited computational resources and helps conducting experiments rapidly. Notably, many augmentation methods [27, 28, 24, 22, 37, 19, 21, 20, 60] have not reported their performance for this popular application. Nevertheless, we evaluate our approach by fine-tuning the baseline model on three datasets Flower102, Aircraft, and Stanford Cars, utilizing ResNet-50 model pre-trained on ImageNet. The results are summarized in Table VI.
@@
On the CUB-Birds-200, GenMix achieved the highest accuracy of $84.45\%$. A similar trend is evident in the Stanford-Cars dataset, where GenMix reached $90.84\%$. GenMix significantly surpassed other augmentation techniques, demonstrating its effectiveness in transfer learning scenarios. As compared to the vanilla baselines, our method achieves absolute performance gains of $2.69\%$, and $1.96\%$ on the used datasets. Since fine-tuning requires substantially fewer computational resources than training from scratch, this experiment highlights the practical importance of GenMix in deep learning applications.
V-E Data Scarcity
@@
Data scarcity presents a significant challenge when training deep neural networks. With only a limited number of examples per class, deep networks may struggle to learn meaningful patterns, which can lead to overfitting and a reduced ability to generalize [84, 85]. To mitigate this, augmentation techniques are commonly applied to increase the amount of training data. In this context, we evaluate the performance of ResNet-18 [66] trained with just $10$ images per class from the original Flower102 dataset. As shown in Table V, GenMix consistently outperforms other mixup approaches, achieving Top-1 validation accuracy is $79.34\%$ and Top-5 accuracy of $76.38\%$. Our method is specifically designed to diversify the training dataset. By using custom conditional prompts, GenMix artificially expands and enriches the dataset, leading to stronger neural network learning.
@@
TABLE VII: Comparison of H-score (%) in the Open Partial Domain Adaptation (OPDA) scenario on the Office-Home, where Clipart (Cl), Art (Ar), Real (Re), and Product (Pr) are different domains. Unified (U) methods are applicable for all scenarios including OPDA, OSDA, and PDA. The SF denotes source data-free, and CF indicates K-means clustering free model adaption. Our method is shown as +GenMix to imply LEAD+GenMix. Boldface represents performance gains of GenMix over LEAD.
Method U SF CF Ar2Cl Ar2Pr Ar2Re Cl2Ar Cl2Pr Cl2Re Pr2Ar Pr2Cl Pr2Re Re2Ar Re2Cl Re2Pr Avg
CMU [86] 56.0 56.9 59.2 67.0 64.3 67.8 54.7 51.1 66.4 68.2 57.9 69.7 61.6
DANCE [87] 61.0 60.4 64.9 65.7 58.8 61.8 73.1 61.2 66.6 67.7 62.4 63.7 63.9
DCC [88] 58.0 54.1 58.0 74.6 70.6 77.5 64.3 73.6 74.9 81.0 75.1 80.4 70.2
OVANet [89] 62.8 75.6 78.6 70.7 68.8 75.0 71.3 58.6 80.5 76.1 64.1 78.9 71.8
GATE [90] 63.8 75.9 81.4 74.0 72.1 79.8 74.7 70.3 82.7 79.1 71.5 81.7 75.6
UniOT [91] 67.3 80.5 86.0 73.5 77.3 84.3 75.5 63.3 86.0 77.8 65.4 81.9 76.6
@@
SHOT-O [92] 32.9 29.5 39.6 56.8 30.1 41.1 54.9 35.4 42.3 58.5 33.5 33.3 40.7
LEAD [54] 62.7 78.1 86.4 70.6 76.3 83.4 75.3 60.6 86.2 75.4 60.7 83.7 75.0
+ GenMix 63.9 83.0 85.9 76.8 78.9 85.3 78.1 62.2 86.9 79.8 62.9 83.0 77.2
TABLE VIII: H-score (%) comparison in the Open-set Domain Adaptation (OSDA) scenario on the Office-Home dataset.
Method U SF CF Ar2Cl Ar2Pr Ar2Re Cl2Ar Cl2Pr Cl2Re Pr2Ar Pr2Cl Pr2Re Re2Ar Re2Cl Re2Pr Avg
CMU [86] 55.0 57.0 59.0 59.3 58.2 60.6 59.2 51.3 61.2 61.9 53.5 55.3 57.6
DANCE [87] 6.5 9.0 9.9 20.4 10.4 9.2 28.4 12.8 12.6 14.2 7.9 13.2 12.9
DCC [88] 56.1 67.5 66.7 49.6 66.5 64.0 55.8 53.0 70.5 61.6 57.2 71.9 61.7
OVANet [89] 58.6 66.3 69.9 62.0 65.2 68.6 59.8 53.4 69.3 68.7 59.6 66.7 64.0
@@
SHOT-O [92] 37.7 41.8 48.4 56.4 39.8 40.9 60.0 41.5 49.7 61.8 41.4 43.6 46.9
LEAD [54] 60.7 70.8 76.5 61.0 68.6 70.8 65.5 59.8 74.2 64.8 57.7 75.8 67.2
+ GenMix 62.8 72.4 76.8 65.1 66.8 71.5 67.3 58.5 74.9 65.7 63.1 79.2 68.7
V-F Domain Adaptation